Comprehensive CVP analysis (LO 1,2,3,5)-rข never unkntand accounting stuer, Blake Dunn yellod, waving the inconw suament he had just received from his accountant in the morning mail. Last month, we sold 1,000 stuffed State Univensity masoots and caned S6850 in operating income. This month, when we sold 1.500, I thought wed make $10 275 But this income statement shows an operating income of $12,100 How can I ever make plans if I cant predict my income? Tm going to give Janice one last chance to explain this to me.he declared as he picked up the phone to call Janice Malley, his accountant Will you try to explain this opcrating income thing to me one more time Blake askod Janice Afer 1saw last mondis income statement.เ ught each mascot wg sold generaod S685.n nct-ncome, now this month, each one gcncrales 58.071 There was no change in the price we paid for cach masoot, so 1 dont understand how this happened. If I had known I was going to have $12,100 in operating income, I would have looked more seriously at adding to our peoduct line Taking a deep breath, Janice replied, Sure, Blake. Id he happy to explain how you made so much more operating income than you were expecting Assume Janices roe. Explain to Blake why his use of operating income per mascot was in o Using the following inome statements, prepare a coetribution margin income statement for Manch February Manch 25,000537 500 Sales revenue Cost of goods sold10000 15000 Gross profa Rent expense Wages expense Shipping expense 250 1875 Ubilitics expense 15.000 22.500 1,500 1500 500 5000 50 750 dvertising espense 750 875 Insurance espense Operating income 6850512.100 00 400 Blake plans to sell 500 stuffed mascots nex month How mach operating income can Hilake expoct to cam next month if he realizes his planned sales? Blake wasnt happy with the projocted income statement you showed him for a sales level of 500 stuffed mascots He wants to know how many stuffied mascots he will nood to sell to cam $3,700 in operating income. As a safety net, he also wants to know how many stuffed mascots he will nood to sell to break even Blake is evaluating two options to increase the mumber of mascots seld nest month First, he believes he can increase sales by advertising in the university newspaper. Blake can purchase a package of 12 ads over the next moeth for a total of S1,.200 He believes the ads will increase the number of stuffed mascots sold from 500 to 960 Asecond option would be to reduce the selling proe Blake believes a10%decrease inthe pr ce will resultin 1,000 mascots sold. Which plan should Blake implement? At what level of sales would he be indifferent between the two plans Just after Blake completed an income projection for 1,200 stuffed mascots, his supplier called to inform him of a X)% increase in cost of goods sold. cflective immediately Blake knowsthat he cannot passthe entire increase on to his customers, but thinks he can pass on half of the 20% increase while suffering only a 5% decrease in units sold. Should Blake respond to the increase in cost of goods sold with an increase in price? Refer back to the original information. Blake has decided to add stadium blankets to his product line. He has found a supplier who will peovide the blankets foe $32, and he plans to sell thom for $55. All other variable costs currently incurred for selling mascots will be incurred for selling blankets at the same rate. Additional fixed costs of $350 per month will be incurred. He believes he can sell one Nanket for every three stuffed mascots How many blankets and stuffed masoots will Blake nced to sell each month in onder to break even
Answer #1

E . Operating profit in case of advertisement

Sales revenue (25x960) 24000
Less : variable cost
- COGS (10000/1000x960) (9600)
- wages (3x960) (2880)
- supplies exp (1250/1000x960) (1200)
Contribution 10320
Less : fixed cost
- rent expense (1500)
- wages (500)
- utilities (750)
- insurance (400)
- advertisement (1200)
Operating profit $5970

Operating profit in case of reduction of sale price

Note : original sale price 25000/1000 = 25

Sale price after discount = 25-10% of 25 = $22.50

Statement showing operating profit

Sales revenue (22.5x1000) 22500
Less : variable cost
- COGS (10000)
- wages (3000)
- shipping exp (1250)
Contribution $8250
Less : fixed cost
- rent exp (1500)
- wages (500)
- utilities (750)
- insurance (400)
Operating profit $5100

Since operating profit is more in case of 1st option, thus, company should advertise by taking pack costing $1200.

F. Cost of goods sold after increase per unit = 10+20% OF 10 = 12.

Out of increase of $2 , he can pass to customer jald burden i.e. he can increase price by $1 . Therefore, new price = $26.

New demand = 1200-5% of 1200 = 1140 units

Statement of operating income

Sales 31200
Less variable cost
COGS (12x1200) (14400)
Wages (3x1200) (3600)
Shipping (1.25x1200) (1500)
Advertisement (0.25x1200) (300)
Contribution $11400
Less : fixed cost
Rent (1500)
Wages (500)
Utilities (750)
Insurance (400)
Advertisement (500)
Operating profit $7750

Yes, it can pass on the burden to customer has it is having positive profit event after passing on burden

G . Contribution on blanket

Selling price 55
Less : cost of purchase (32)
Salary (3)
Shipping exp (1.25)
Advertisement exp (0.25)
Contribution $18.5

Weighted average contribution = 10.5x3+18.5x1/4 = $12.5

Total new fixed cost as per original plus new = 3650+350 = $4000

Weighted average break even = 4000/12.5 = 320 units

Sale of blanket = 320/4 = 80 units

Sale of stuffed mascots = 320x3/4 = 240 units

Note : This solution was in continuation of precious solution that I have given to you in that I have solved first four parts and due lack of time I wasn't able that time to solve last three . Thst i have solved now for you. Thanks